A rectangle has a length of 4x-10 and a width of 8x-2. If the perimeter is 240 inches what are the length and width?

Answer by Boreal(15235)

Half perimeter = 120 inches=L+W=4x-10+8x-2=12x-12
12x=132
x=11
width is 86 inches
length is 34 inches
They add to 120 inches